The Condé Nast Traveler Gold List provides ratings for the top 20 small cruise ships. The data shown below are the scores each ship received based upon the results from Condé Nast Traveler's Annual Readers' Choice Survey. Each score represents the percentage of respondents who rated a ship as excellent or very good on several criteria, including Shore Excursions and Food/Dining. An overall score was also reported and used to rank the ships. The highest ranked ship, the Seabourn Odyssey, has an overall score of 94.4, the highest component of which is 97.8 for Food/Dining.

Ship Overall Shore
Excursions
Food/Dining
Seabourn Odyssey 94.4 90.9 97.8
Seabourn Pride 93.0 84.2 96.7
National Geographic Endeavor 92.9 100.0 88.5
Seabourn Sojourn 91.3 94.8 97.1
Paul Gauguin 90.5 87.9 91.2
Seabourn Legend 90.3 82.1 98.8
Seabourn Spirit 90.2 86.3 92.0
Silver Explorer 89.9 92.6 88.9
Silver Spirit 89.4 85.9 90.8
Seven Seas Navigator 89.2 83.3 90.5
Silver Whisperer 89.2 82.0 88.6
National Geographic Explorer 89.1 93.1 89.7
Silver Cloud 88.7 78.3 91.3
Celebrity Xpedition 87.2 91.7 73.6
Silver Shadow 87.2 75.0 89.7
Silver Wind 86.6 78.1 91.6
SeaDream II 86.2 77.4 90.9
Wind Star 86.1 76.5 91.5
Wind Surf 86.1 72.3 89.3
Wind Spirit 85.2 77.4 91.9

(a) Determine an estimated regression equation that can be used to predict the overall score given the score for Shore Excursions. (Round your numerical values to two decimal places. Let x1 represent the Shore Excursions score and y represent the overall score.)

ŷ =

(b) Consider the addition of the independent variable Food/Dining. Develop the estimated regression equation that can be used to predict the overall score given the scores for Shore Excursions and Food/Dining. (Round your numerical values to two decimal places. Let x1 represent the Shore Excursions score, x2 represent the Food/Dining score, and y represent the overall score.)

ŷ =

(c) Predict the overall score for a cruise ship with a Shore Excursions score of 78 and a Food/Dining Score of 92. (Round your answer to one decimal place.)
Answer #1

We input the data set in Excel and use the "Regression" option under Data > Data Analysis to build two regression models - the first with y on x1 and second with y on x1 and x2. The screenshot of the data set and the outputs is shown below.
B С D E F H - J K N 0 0 P R S M SUMMARY OUTPUT SUMMARY OUTPUT 1 ly 2 A x1 94.4 93 92.9 x2 90.9 3 84.2 97.8 96.7 88.5 4 100 5

(a) The regression line is: \hat{y} = 69.30 + (0.23) x_1 .
(b) The regression line is: \hat{y} = 45.18 + (0.25) x_1 + (0.25) 12 .
(c) Using the regression line in part (b) and the values of x1 = 78 and x2 = 92, we get predicted overall score = 45.18 + (0.25 * 78) + (0.25 * 92) = 87.7.